# Npix2Cpix: A GAN-based Image-to-Image Translation Network with Retrieval-Classification Integration for Watermark Retrieval from Historical Document Images

### v1

@@ III-B Image to Image Translation: Npix2Cpix
To eliminate the handwriting from the input images as well as to remove the watermark background’s discontinuity and clutter, we propose a conditional GAN-based technique for the first step of our pipeline. To put it in another way, this proposed network will receive a raw noisy watermarked image, as input and will generate a noise-free image, which is the generated watermarked image, with only the watermark. From the literature study, it is found that GANs are one of the best ways to generate a clean image from a raw noisy watermarked image, but surprisingly has not yet been enough explored for the historical watermark retrieval domain. This application of GANs is vastly different from the traditional uses of GANs as well as previous watermark retrieval techniques employed in this domain. Usually, GANs are used to generate new, synthetic instances of images that can pass for real data, whereas the scope of GAN usage in this domain is limited to deblurring or removing compression artifacts. Using a conditional adversarial network, [44] shows that it is efficient at a number of tasks, including colorizing pictures, reconstructing objects from edge maps, and synthesizing images from label maps.
@@
In order to train the discriminator, firstly score for patches from the pair of noisy watermarked images, , and ground truth image, is calculated. Then prediction of patches from another pair containing the noisy watermarked image, , and the generated watermarked image, is obtained as shown in Fig. 5.
for number of training iterations do
@@
The discriminator loss, is the combination of two different losses, mean square error (MSE) [46] calculated from, and calculated from their corresponding ground truths. After obtaining patch-wise predictions from the discriminator model, Discriminator generated loss, and Discriminator real loss, calculated for both pairs. As mentioned before, the goal of the discriminator is to discriminate between real and generated images and predict correctly. The intuition behind these losses is that, optimizing will teach the discriminator to recognize real images and predict 1 for them, whereas optimizing will teach the discriminator to recognize generated images and predict 0 for them the total discriminator loss, is the mean of and . So, the discriminator loss can be written as eqn. 1
The main target of the generator is to reconstruct the noisy watermarked image, in such a way that the generated watermarked image, is close enough to the ground truth image, , as well as confuse the discriminator. That is why, here the generator loss, is calculated from MSE of discriminator’s output from the pair of and against and L1 loss [47] calculated between and . These losses, denoted by and , respectively, are combined to obtain . Here, fulfills the goal of the generator by directly comparing the generated watermarked image, with ground truth image, and by confusing the discriminator through optimizing in the opposite direction of the gradient updated at discriminator training. Here, gradients are updated in the opposite direction of . So, the generator loss can be described as eqn. 2
In order to improve the denoising performance of the proposed Npix2Cpix GAN, it is important to concurrently minimize both and by optimizing the network parameters. Any imbalance in training the generator and discriminator sometimes leads the training process to diverge or collapse. Upon successful training, the generator of the trained GAN is expected to produce a denoised watermarked image from a degraded historical watermarked image. This denoised watermarked image is then utilized in algorithm 2 for classifying the exact class of watermarks which is explained elaborately in section III-C.
III-C Metric-Based Classifier: Siamese Network
@@
In this paper, the challenges associated with classifying noisy watermarks are clearly outlined in subsection III-A, which are comprehensively addressed through the proposed conditional U-Net-based GAN. However, there may still arise some challenging scenarios where documents remain significantly degraded and may not be fully denoised. As a result, efforts are focused on improving the feature extraction functionality of the Siamese network so that important features can be extracted from marginally denoised watermarked images. In order to accomplish this, adjustments are implemented to the conventional Siamese network, enabling the smooth incorporation of diverse pretrained networks such as ResNet and EfficientNet for the purpose of extracting meaningful features. In addition, to enhance the discriminatory capabilities of the network, the cosine distance is utilized to compute the similarity between image pairs, instead of the conventional Euclidean distance.
for number of training iterations do
@@
The architecture of our implementation for siamese-based one-shot classification consists of three components: a feature extractor, a feature comparison metric, and a suitable loss function. At first, image features and are extracted from input images and respectively after the images are cleaned with NPix2CPix GAN. The methodology for the proposed conditional GAN-based denoising is covered in detail in subsection III-B. The distance, between these two features, is calculated using the distance metric as shown in Fig 6. The employed unit of measurement is the cosine similarity [48]. The ground truth label 0 is used if the image pair is taken from the same class and a label of 1 is given if images from the pair are from different classes, which is denoted by . When trained with these labels, the model gains the ability to distinguish between features belonging to different classes and to concentrate on watermark similarities.
To train the model for detecting the similarity, we need a loss function that can work with the distance, calculated from extracted features. The loss function has to minimize of the input pair from the same class and maximize up to a margin, if the input images are from a different class, as depicted in Algorithm 2. The Contrastive Loss, a differentiable loss function, is used as Siamese Loss, as it can fulfill this requirement. Our implementation of the loss function, is denoted by eqn. 3
The final prediction of the model is represented here by , the cosine distance () between two extracted features. The dataset label is denoted by the value , and the margin value, represents the expected variation in pair distance scores between similar and dissimilar pairs.
@@
III-C1 Training Procedure of Siamese Network
First, our suggested U-net-based conditional GAN is used to denoise all the impaired images from 100 classes (50 images/class). The cleaned images of 100 classes are then split into a 60:40 ratio for the training and inference procedure. During the training phase, the model is trained using a similarity learning strategy. In order to do this, it is necessary to use image pairings from similar and dissimilar classes. These pairs are produced using the cleaned 60% (60 classes) data of ’A Classification’ dataset. A randomly selected class from the train portion data is picked, and two randomly selected images from that class are taken to create similar pairings. Again two classes are arbitrarily selected for the dissimilar pairs, and then one image from each class is selected for the pair. 1000 additional image pairings are used for validation after each 10,000 training pairs for each epoch.
@@ III-C1 Training Procedure of Siamese Network
III-C2 Inference Procedure of Siamese Network
@@ III-C3 One-shot classification
For the purpose of classification, one-shot classification with a Siamese network is used. The difficult issue of historical watermark categorization is addressed using a one-shot learning-based classification technique. The one-shot test can show that the proposed GAN-based technique can extract the clean watermark in the majority of situations when genuine clean images are used as the support set and produced clean images are used as the target set for classification. To distinguish watermarks from dissimilar classes in the compressed feature space, the suggested architecture is built on a Siamese network with several backbones (Resnet, Mobilenet, and Efficientnet).
III-D An Additional Experiment for the Proposed U-net
@@
Next, the performance of the proposed U-net-based design for the generator model is investigated. In Fig. 8, with the help of a simplified block diagram, the task of extracting clean watermarks from the synthetically generated documents is explained using the proposed U-net-based GAN generator. Synthetic documents that have been degraded by the addition of random handwriting are used to estimate how well the suggested model performs. Moreover, random Gaussian noise is also incorporated into the document, as shown in Fig. 7, to make the task more challenging.
@@
Fig. 8 depicts the generated document, , which is quite similar to the ground truth, , where, is the synthetically generated document with handwriting and noise. This experiment is carried out to determine whether the proposed U-net-based generator is robust enough to denoise the document which belongs to a different domain from the noise-corrupted synthetic document and retrieve the watermark.
IV Result and Analysis
@@
In this section, experimental details of the proposed work and comparative performance analysis with existing methods are presented. The training of the proposed GAN-based model is conducted on Nvidia P100 GPU, requiring approximately 10 hours to complete. Subsequently, for the Siamese-based one-shot classification task, the same Nvidia P100 GPU is employed, with a training duration of 4.5 hours. It is important to note that the time required for specific tasks may vary depending on the number of classes involved. In the context of this study, the reported time duration pertains to experiments conducted with the dataset comprising 100 classes.
@@
A few metrics [49], that are used to evaluate how precisely the watermark is extracted, are demonstrated to prove the viability of the suggested technique. In Fig. 9, a generated image from the noisy sample, which is quite similar to the ground truth, is shown. Here extracted watermark is marked by a green box for better visualization. The table contains a few parametric metrics such as MSE, PSNR, SSIM, and BRISQUE.
@@
Mean squared error (MSE) is the most widely used estimator of an image quality metric. Lower values of MSE are better. MSE is described by eqn. 4.
(4) - PSNR:
The peak signal-to-noise ratio (PSNR) between two images is used to compare the original picture quality to the transformed image quality. The PSNR increases with the quality of the compressed or reconstructed image. PSNR and MSE are used to gauge the quality of image compression. The mathematical expression of PSNR is shown in eqn. 5
@@
In order to measure the similarity between two digital images or videos, the structural similarity index metric (SSIM) is used [50]. An original, uncompressed, distortion-free reference image is used to evaluate the quality of the test image. The SSIM index can provide precise information on how closely the generated watermark is to its ground truth. Between two images and , the SSIM can be expressed as eqn. 6
@@ IV-A Qualitative evaluation
IV-B Quantitative Evaluation
For the quantitative assessment of the performance of the proposed watermarked image denoising method, initially, a few image quality metrics are computed for both actual noisy images as well as generated images, such as MSE, RMSE, SSIM, and BRISQUE. Results are compared in Table I. From Table I, it is quite evident that all quantitative measures are improved when generated images are compared to the actual noisy images.
@@ IV-B Quantitative Evaluation
The results on the selected metrics, in the case of synthetic data, are shown in Table II. Similar performance is also observed here, which provides additional support for the dependability of the proposed architecture in a different domain.
@@
The BRISQUE score comparison is shown with several images from a definite class in Table III. Typically, the BRISQUE score spans between 0 and 100. The improved perceptual quality of the image is reflected in lower score values. Table III shows that the generated images using the proposed technique have a score that is rather near to the ground truth images in both domains.
@@
| Ground Truth | 23.324 |
| Noisy Image | 30.518 |
| Generated Image | 24.139 |
| Ground Truth (Syn.) | 56.271 |
| Noisy Image (Syn.) | 88.663 |
| Generated Clean (Syn.) | 54.033 |
For the purpose of classifying the extracted watermarks, a Siamese-based one-shot method is demonstrated here. As previously discussed, the feature extractor is constructed in such a manner that the backbone network can be simply altered for the experiment of the one-shot method. Results using several backbones are shown in Table IV, which shows the classification accuracy in the one-shot approach. The findings for the feature extractors Resnet18 [52], Efficientnetb0 [53], and MobilenetV3 (small) [54] are presented in Table IV, with Resnet18 architecture providing the best one-shot test accuracy (95%). Due to the extensive use of parameters in this design, Resnet18 is employed here, which outperforms the other architectures. Efficientnet is employed because of its ability to optimize both parameters and performance. Finally, one-shot performance with fewer parameters is checked using the MobilenetV3 (small) results.
@@
The watermark classification accuracy achieved through our proposed method is compared with some recent methods dealing with the same problem and results are shown in Table V.
It is observed from Table V that, the proposed method achieves 95% classification accuracy, which is much higher than the accuracies obtained by some existing works such as the local spatially-aware approach and Siamese network with thresholding-based enhancement. Table V clearly shows that the Siamese network with the proposed U-Net-based conditional GAN performs better than the recent best methods. In particular, performance is noticeably degraded when this proposed GAN-based denoising method has not been utilized. This is because the feature extractor has to work with historical watermarked images that are severely distorted to classify them. However, using the proposed method to remove noise from the historical watermarked document leads to a major boost in classification accuracy, since the feature extractor now has to handle a noise-free watermarked image. So, it can be concluded that the proposed one-shot classification with a GAN-based pre-processing is more efficient than the existing state-of-the-art method.
@@
For experimentation with the proposed one-shot classification method images are resized to 5 scales. In Table VI, results of classification accuracies obtained for several scales are presented. Scale-i corresponds to the specific size , , , , , respectively for i = 1,2,…,5. In all the scales, the Resnet18 backbone provides consistently better classification accuracy and exhibits relatively less variation with respect to scales.
V Conclusion
This research demonstrates how, despite the fact that watermarks are represented in a variety of ways, deep learning-based algorithms are highly useful for identifying and extracting watermarks from historical document images. This work proposes a U-net-based conditional GAN that can clean noisy historical watermarked documents and retrieve the specific watermarked image. The main difference between the proposed studies and existing ones is the utilization of a generative adversarial network (GAN)-based image-to-image translation method for watermark extraction, followed by a Siamese-based one-shot classification approach, which overcomes issues associated with traditional denoising and classification methods while addressing the scarcity of data. The proposed network is further tested using images from a different domain, and in this instance, rich-quality watermarks are also retrieved. Several metrics and one-shot classification accuracy are used to evaluate the quality of extracted watermarks. Extensive experiments show that the proposed technique outperforms some recent methods when it comes to classifying the denoised watermark utilizing a one-shot approach.

### v3

@@ 3.2 Image to Image Translation: Npix2Cpix
To eliminate the handwriting from the input images as well as to remove the watermark background’s discontinuity and clutter, we propose a conditional GAN-based technique for the first step of our pipeline. To put it in another way, this proposed network will receive a raw noisy watermarked image, as input and will generate a noise-free image, which is the generated watermarked image, with only the watermark. From the literature study, it is found that GANs are one of the best ways to generate a clean image from a raw noisy watermarked image. Still, it surprisingly has not yet been explored enough for the historical watermark retrieval domain. This application of GANs is vastly different from the traditional uses of GANs as well as previous watermark retrieval techniques employed in this domain. Usually, GANs are used to generate new, synthetic instances of images that can pass for real data, whereas the scope of GAN usage in this domain is limited to deblurring or removing compression artifacts. Using a conditional adversarial network, [46] shows that it is efficient at a number of tasks, including colorizing pictures, reconstructing objects from edge maps, and synthesizing images from label maps.
@@
In order to train the discriminator, firstly score for patches from the pair of noisy watermarked images, , and ground truth image, is calculated. Then prediction of patches from another pair containing the noisy watermarked image, , and the generated watermarked image, is obtained as shown in Fig. 3 and Fig. 5.
for the number of training iterations do
@@
The discriminator loss, is the combination of two different losses, mean square error (MSE) [48] calculated from, and calculated from their corresponding ground truths. After obtaining patch-wise predictions from the discriminator model, discriminator generated loss, and discriminator real loss, calculated for both pairs. As mentioned before, the goal of the discriminator is to discriminate between real and generated images and predict correctly. The intuition behind these losses is that optimizing will teach the discriminator to recognize real images and predict 1 for them, whereas optimizing will teach the discriminator to recognize generated images and predict 0 for them. The total discriminator loss, is the weighted mean of and . So, the discriminator loss can be written as Eq. (1).
Here, is the discriminator generated loss weight and is the discriminator real loss weight.
The main target of the generator is to reconstruct the noisy watermarked image, in such a way that the generated watermarked image, is close enough to the ground truth image, , as well as confuse the discriminator. That is why, here the generator loss, is calculated from MSE of discriminator’s output from the pair of and against and L1 loss [49] calculated between and . These losses, denoted by and , respectively, are combined to obtain . Here, fulfills the goal of the generator by directly comparing the generated watermarked image, with ground truth image, and does it by confusing the discriminator through optimizing in the opposite direction of the gradient updated at discriminator training. Here, gradients are updated in the opposite direction of . So, the generator loss, can be described as Eq. (2)
Here, is the generator’s generated loss weight, and is the generator’s real loss weight.
In our approach to denoising noisy historical documents using the proposed Npix2Cpix GAN, we have chosen to keep the weights for both real and generated losses equal. This decision is based on the need to maintain balance in GAN training. By keeping the weights equal, we ensure that both real and generated losses contribute equally, preventing either the generator or the discriminator from becoming too dominant. This balance is crucial for achieving stable and effective training dynamics.
In order to improve the denoising performance of the proposed Npix2Cpix GAN, it is important to concurrently minimize both and by optimizing the network parameters. Any imbalance in training the generator and discriminator sometimes leads the training process to diverge or collapse. Upon successful training, the generator of the trained GAN is expected to produce a denoised watermarked image from a degraded historical watermarked image. This denoised watermarked image is then utilized in Algorithm 2 for classifying the exact class of watermarks which is explained elaborately in section 3.3.
3.3 Metric-Based Classifier: Siamese Network
@@
In this paper, the challenges associated with classifying noisy watermarks are clearly outlined in subsection 3.1, which are comprehensively addressed through the proposed conditional U-Net-based GAN. However, there may still arise some challenging scenarios where documents remain significantly degraded and may not be fully denoised. As a result, efforts are focused on improving the feature extraction functionality of the Siamese network so that important features can be extracted from marginally denoised watermarked images. In order to accomplish this, adjustments are implemented to the conventional Siamese network, enabling the smooth incorporation of diverse pretrained networks such as ResNet, MobileNet, and EfficientNet for the purpose of extracting meaningful features. In addition, to enhance the discriminatory capabilities of the network, the cosine distance is utilized to compute the similarity between image pairs, instead of the conventional Euclidean distance.
for the number of training iterations do
@@
The architecture of our implementation for Siamese-based one-shot classification consists of three components: a feature extractor, a feature comparison metric, and a suitable loss function. At first, image features and are extracted from input images and respectively after the images are cleaned with the proposed Npix2Cpix GAN. The methodology for the proposed conditional GAN-based denoising is covered in detail in subsection 3.2. The distance, between these two features, is calculated using the distance metric as shown in Fig 6. The employed unit of measurement is the cosine similarity [50]. The ground truth label 0 is used if the image pair is taken from the same class and a label of 1 is given if images from the pair are from different classes, which is denoted by . When trained with these labels, the model gains the ability to distinguish between features belonging to different classes and to concentrate on watermark similarities.
To train the model for detecting the similarity, we need a loss function that can work with the distance, calculated from extracted features. The loss function has to minimize of the input pair from the same class and maximize up to a margin, if the input images are from a different class, as depicted in Algorithm 2. The Contrastive loss, a differentiable loss function, is used as Siamese loss, as it can fulfill this requirement. Our implementation of the loss function, is denoted by Eq. (3)
The final prediction of the model is represented here by , the cosine distance between two extracted features. The dataset label is denoted by the value , and the margin value, represents the expected variation in pair distance scores between similar and dissimilar pairs.
@@
3.3.1 Training Procedure of Siamese Network
First, our suggested U-net-based conditional GAN is used to denoise all the impaired images from 100 classes (50 images/class). The cleaned images of 100 classes are then split into a 60:40 ratio for the training and inference procedure. During the training phase, the model is trained using a similarity learning strategy. In order to do this, it is necessary to use image pairings from similar and dissimilar classes. These pairs are produced using the cleaned 60% (60 classes) data of the ’A Classification’ dataset. A randomly selected class from the train portion data is picked, and two randomly selected images from that class are taken to create similar pairings. Again two classes are arbitrarily selected for the dissimilar pairs, and then one image from each class is selected for the pair. 1000 additional image pairings are used for validation after each 10,000 training pairs for each epoch.
@@ 3.3.1 Training Procedure of Siamese Network
3.3.2 Inference Procedure of Siamese Network
@@ 3.3.3 One-shot classification
For the purpose of classification, one-shot classification with a Siamese network is used. The difficult issue of historical watermark categorization is addressed using a one-shot learning-based classification technique. The one-shot test shows that the proposed GAN-based technique can denoise the watermarked documents in the majority of situations when genuine clean watermarked images are used as the support set and generated watermarked images are used as the target set for classification. To distinguish watermarks from dissimilar classes in the compressed feature space, the suggested architecture is built on a Siamese network with several backbones (ResNet, MobileNet, and EfficientNet).
3.4 An Additional Experiment for U-net
@@
Next, the performance of the proposed U-net-based design for the generator model is investigated. In Fig. 8, with the help of a simplified block diagram, the task of extracting clean watermarks from the synthetically generated documents is explained using the generator of the proposed U-net-based GAN. Synthetic documents that have been degraded by the addition of random handwriting are used to estimate how well the suggested model performs. Moreover, random Gaussian noise is also incorporated into the document, as shown in Fig. 7, to make the task more challenging.
@@
Fig. 8 depicts the generated watermarked synthetic image, , which is quite similar to the ground truth image, , where, is the synthetic watermarked image with handwriting and noise. This experiment is carried out to test whether the proposed U-net-based generator is robust enough to denoise the document which belongs to a different domain from the noise-corrupted synthetic document and retrieve the watermark.
4 Result and Analysis
@@
In this section, experimental details of the proposed work and comparative performance analysis with existing methods are presented. The training of the proposed GAN-based model is conducted on a Tesla P100 GPU, requiring approximately 10 hours to complete. Subsequently, for the Siamese-based one-shot classification task, the same Tesla P100 GPU is employed, with a training duration of 4.5 hours. The experimental setup utilized a single Tesla P100 GPU with 3584 CUDA cores and 16GB GDDR6 VRAM, a single-core hyper-threaded Xeon processor at 2.2GHz, 15.26GB of RAM, and 155GB of available disk space. It is important to note that the time required for specific tasks may vary depending on the number of classes involved. In the context of this study, the reported time duration pertains to experiments conducted with the dataset comprising 100 classes. Additionally, the training configurations and hyperparameters used for training the proposed U-net-based GAN and Siamese network are described in Table 1.
| Hyperparameters | Npix2Cpix | Siamese network |
| Batch size | 16 | 8 |
| Learning rate | 0.0002 | 0.001 |
| Betas | [0.5, 0.999] | [0.9, 0.999] |
| Optimizer | Adam | Adam |
| Epochs | 200 | 35 |
| Image size |
@@
A few metrics [51], that are used to evaluate how precisely the watermark is extracted, are demonstrated to prove the viability of the suggested technique. In Fig. 9, a generated image from the noisy sample, which is quite similar to the ground truth, is shown. Here, the extracted watermark is marked by a green box for better visualization. The table contains a few parametric metrics such as MSE, PSNR, SSIM, and BRISQUE.
@@
Mean squared error (MSE) is the most widely used estimator of an image quality metric. Lower values of MSE are better. MSE is described by Eq. (4).
(4) where is the ground truth image and is the generated image.
- PSNR:
The peak signal-to-noise ratio (PSNR) between two images is used to compare the original picture quality to the transformed image quality. The PSNR increases with the quality of the compressed or reconstructed image. PSNR and MSE are used to measure the quality of image compression. The mathematical expression of PSNR is shown in Eq. (5).
@@
In order to measure the similarity between two digital images or videos, the structural similarity index metric (SSIM) is used [52]. An original, uncompressed, distortion-free reference image is used to evaluate the quality of the test image. The SSIM index can provide precise information on how closely the generated watermark is to its ground truth. Between two images and , the SSIM can be expressed as Eq. (6).
@@ 4.1 Qualitative evaluation
4.2 Quantitative Evaluation
For the quantitative assessment of the performance of the proposed watermarked image denoising method, initially, a few image quality metrics are computed for both actual noisy images as well as generated images, such as MSE, RMSE, SSIM, and BRISQUE. Results are compared in Table 2. From Table 2, it is quite evident that all quantitative measures are improved when generated images are compared to the actual noisy images.
@@ 4.2 Quantitative Evaluation
The results on the selected metrics, in the case of synthetic data, are shown in Table 3. Similar performance is also observed here, which provides additional support for the dependability of the proposed architecture in a different domain.
@@
The BRISQUE score comparison is shown with several images from a definite class in Table 4. Typically, the BRISQUE score spans between 0 and 100. The improved perceptual quality of the image is reflected in lower score values. Table 4 shows that the generated images using the proposed technique have a score that is rather near to the ground truth images in both domains.
@@
| Ground Truth Image | 23.324 |
| Noisy Watermarked Image | 30.518 |
| Generated Watermarked Image | 24.139 |
| Ground Truth Image (Syn.) | 56.271 |
| Noisy Watermarked Image (Syn.) | 88.663 |
| Generated Watermarked Image (Syn.) | 54.033 |
The watermark classification accuracy achieved through our proposed method is compared with some recent methods dealing with the same problem and results are shown in Table 5.
It is observed from Table 5 that, the proposed method achieves 95% classification accuracy, which is much higher than the accuracies obtained by some existing works such as the local spatially-aware approach and Siamese network with thresholding-based enhancement. So, it can be concluded that the proposed one-shot classification with a GAN-based pre-processing is more efficient than the existing state-of-the-art methods.
4.3 Ablation Experiment
4.3.1 Effect of Npix2Cpix GAN
To show the effectiveness of the proposed GAN-based denoising, we conducted experiments to compare the performance of the one-shot approach with and without GAN-based denoising prior to the classification stage. The results of these experiments are detailed and compared in Table 6.
| Method | Accuracy |
| Siamese Network | 76.2% |
| Proposed Npix2Cpix + Siamese Network | 95% |
Table 6 clearly shows that the Siamese network with the proposed U-Net-based conditional GAN performs better. In particular, performance is noticeably degraded when the proposed GAN-based denoising method has not been utilized. This is because the feature extractor of the Siamese network has to work with historical watermarked images that are severely distorted for classification. However, using the proposed method to remove noise from the historical watermarked document leads to a major boost in classification accuracy, since the feature extractor now has to handle a noise-free watermarked image.
4.3.2 Effect of different feature extractor of Siamese network
For the purpose of classifying the denoised watermarked image, a Siamese-based one-shot method is demonstrated in this work. As previously discussed, the feature extractor is constructed in such a manner that the backbone network can be simply altered for the one-shot method experiment. Results using several backbones are shown in Table 7, which shows the classification accuracy in the one-shot approach. The findings for the feature extractors such as Resnet18 [54], Efficientnetb0 [55], and MobilenetV3 (small) [56] are presented in Table 7, with Resnet18 architecture providing the best one-shot test accuracy (95%). Due to the extensive use of parameters in this design, Resnet18 is employed here, which outperforms the other architectures. Efficientnet is employed because of its ability to optimize both parameters and performance. Finally, one-shot performance with fewer parameters is checked using the MobilenetV3 (small) results.
@@ 4.3.2 Effect of different feature extractor of Siamese network
4.3.3 Effect of image shape variations
For experimentation with the proposed one-shot classification method images are resized to 5 scales. In Table 8, results of classification accuracies obtained for several scales are presented. Scale-i corresponds to the specific size , , , , , respectively for i = 1,…,5. In all the scales, the Resnet18 backbone provides consistently better classification accuracy and exhibits relatively less variation with respect to scales.
5 Limitations and Future Work
This paper presents a modified U-net-based conditional generative adversarial image-to-image translation network for watermark retrieval from historical documents and a Siamese-based one-shot approach for classifying the retrieved watermarks. The proposed method is validated using a large-scale dataset containing ancient watermarked images. While the method shows promise for denoising noisy watermarked documents, there are a few limitations. Historical documents often exhibit varying degrees of noise, including severe degradation and complex clutter, which may impact the model’s performance. In extreme cases, the model may struggle to effectively denoise images, potentially leading to incomplete restoration of watermarks. Additionally, the diversity and complexity of watermarks pose a challenge to the model’s scalability. Future research should focus on developing more robust denoising techniques incorporating additional pre-processing steps to handle variations in noise levels. Expanding the training dataset to include more diverse watermark samples from different domains and employing advanced feature extraction methods could further enhance the model’s effectiveness.
6 Conclusion
This research demonstrates how, despite the fact that watermarks are represented in a variety of ways, deep learning-based algorithms are highly useful for identifying and extracting watermarks from historical document images. This work proposes a U-net-based conditional GAN that can clean noisy historical watermarked documents and retrieve the specific watermarked image. The main difference between the proposed studies and existing ones is the utilization of a generative adversarial network (GAN)-based image-to-image translation method for watermark extraction, followed by a Siamese-based one-shot classification approach, which overcomes issues associated with traditional denoising and classification methods while addressing the scarcity of data. The proposed network is further tested using images from a different domain, and in this instance, rich-quality watermarks are also retrieved. Several metrics and one-shot classification accuracy are used to evaluate the quality of denoised watermarked images. Extensive experiments show that the proposed technique outperforms some recent methods when it comes to classifying the denoised watermarked image utilizing a one-shot approach.